Electronic Journal: Southwest Journal of Pure and Applied Mathematics Internet: http://rattler.cameron.edu/swjpam.html ISSN 1083-0464 Vol. No. 02, 1996, pp. 1–5 Submitted: February 3, 1996. Published: July 8, 1996.

PREDICTIVE DISTRIBUTION FOR POLYTOMOUS LOGISTIC MODELS

WITH NONLINEAR LINK FUNCTIONS

M.A. Tabatabai and K.P. Singh

Abstract. This article is concerned with predictive distribution for polytomous logistic models with nonlinear link functions. Within a Bayesian framework, an approximation of the predictive distribution for determining probabilities is presented. We describe this method using nonlinear link functions. The generalized linear link function is simply a special case of nonlinear models. All of the generalized logistic models require nonlinear estimation procedures to evaluate maximum likelihood estimates.

(1991) A.M.S. Subject Classification Codes. 34.

Key Words and Phrases. polytomous logistic model,link function,predictive distribution

Cameron University, Lawton, Oklahoma E-Mail Address: mohammat@cameron.edu University of Alabama, Birmingham, Alabama \begin{matrix}c \\ ◯ \end{matrix}1996 Cameron University

Typeset by AMS-TEX